# Diffusion In Diffusion: Breaking the Autoregressive Bottleneck in Block Diffusion Models

### v1

@@ Abstract
Block diffusion language models, operating as semi-autoregressive paradigms, combine the strengths of both autoregressive and diffusion paradigms. However, their strict unidirectional block dependencies introduce irreversibility and sacrifice the global planning capabilities for which diffusion models are renowned. In order to address these issues, we propose Diffusion in Diffusion — a ’draft-then-refine’ framework designed to overcome the irreversibility and myopia problems inherent in block diffusion models. Our approach first employs block diffusion to generate rapid drafts using small blocks, then refines these drafts through global bidirectional diffusion with a larger bidirectional receptive field. We utilise snapshot confidence remasking to identify the most critical tokens that require modification, and apply mix-scale training to expand the block diffusion model’s global capabilities. Empirical results demonstrate that our approach sets a new benchmark for discrete diffusion models on the OpenWebText dataset. Using just 26% of the fine-tuning budget of baseline models, we reduce generative perplexity from 25.7 to 21.9, significantly narrowing the performance gap with autoregressive models.
@@ 1 Introduction
Today, language modeling primarily follows two mainstream paradigms: autoregressive (AR) models and discrete diffusion models. AR models (Vaswani et al., 2017; Brown et al., 2020) generate tokens sequentially from left to right and achieve efficient inference through key-value (KV) caching during inference. However, they suffer from linearly increasing inference latency. Conversely, discrete diffusion models (Austin et al., 2021; Lou et al., 2024) enable parallel generation with flexible control over efficiency and quality. Yet, the lack of KV caching hinders efficient inference, and they often underperform AR models on perplexity metrics. Recently, block-diffusion language models (BD3-LMs) (Arriola et al., 2025) have emerged as a semi-AR architecture. By executing diffusion in parallel within “blocks” while employing autoregressive sequence generation between blocks, BD3-LMs successfully bridge the gap between the two paradigms: achieving state-of-the-art performance in non-autoregressive methods while effectively leveraging KV caching for inference acceleration.
However, the standard block diffusion model inevitably introduces myopia and irreversibility due to the inclusion of autoregressive components: the model can only access bidirectional contextual information within a single block, while strictly adhering to an autoregressive causal paradigm between blocks. Furthermore, once diffusion within a block is complete, that block becomes “frozen” and permanently serves as an unupdatable “reference-only” contextual block during subsequent block processing. This leads to the same error accumulation problem as AR models. Early generated tokens, even if erroneous or locally inconsistent, become locked and immutable. Consequently, subsequent token generation is affected, preventing the model from correcting these issues. Therefore, due to the lack of exposure to the global context during generation, BD3-LM struggles to maintain long-range consistency, resulting in a partial loss of global planning capability.
@@
In this study, we propose a novel generative framework called Diffusion in Diffusion (also known as Structured Block Diffusion), designing a multi-stage “draft-then-revise” paradigm to overcome the aforementioned auto-regressive global information bottleneck. Inspired by the human writing process—rapidly drafting a first version followed by holistic revision to ensure coherence—our approach decomposes generation into progressive stages: During the initial drafting phase, the model employs a small blocksize for fast, efficient generation, focusing first on specific grammar and local content. In subsequent revision phases, the model re-examines and refines the previously generated sequence using a larger blocksize (global receptive field). We also innovatively introduce a Remasking mechanism into Block-Diffusion Models, which selectively transforms tokens within the generated sequence back into noise. This enables the model to resample these positions within a more holistic context for optimization.

### v2

# Diffusion In Diffusion: Reclaiming Global Coherence in Semi-Autoregressive Diffusion

@@ Abstract
One of the most compelling features of global discrete diffusion language models is their global bidirectional contextual capability. However, existing block-based diffusion studies tend to introduce autoregressive priors, which, while offering benefits, can cause models to lose this global coherence at the macro level. To regain global contextual understanding while preserving the advantages of the semi-autoregressive paradigm, we propose Diffusion in Diffusion — a ’draft-then-refine’ framework designed to overcome the irreversibility and myopia problems inherent in block diffusion models. Our approach first employs block diffusion to generate rapid drafts using small blocks, then refines these drafts through global bidirectional diffusion with a larger bidirectional receptive field. We utilize snapshot confidence remasking to identify the most critical tokens that require modification, and apply mix-scale training to expand the block diffusion model’s global capabilities. Empirical results demonstrate that our approach sets a new benchmark for discrete diffusion models on the OpenWebText dataset. Using only 26% of the fine-tuning budget of baseline models, we reduce generative perplexity from 25.7 to 21.9, significantly narrowing the performance gap with autoregressive models.
@@ 1 Introduction
Today, research on discrete diffusion models (Austin et al., 2021; Lou et al., 2024) is divided into two major paradigms: global diffusion models and block diffusion models (also known as Semi-Autoregressive diffusion models). Global diffusion models (Sahoo et al., 2024; Nie et al., ) possess strong global planning capabilities and complete bidirectional context, resulting in text with high global consistency. However, they also prevent the application of Key-Value Cache—the foundation for inference acceleration in modern language models (Vaswani et al., 2017; Brown et al., 2020) —leading to an $O(L^{2})$ complexity during inference and extremely slow processing. In contrast, block diffusion models (Arriola et al., 2025) represent an emerging semi-AR architecture that incorporates autoregressive priors. They enable parallel diffusion execution within “blocks” while employing inter-block autoregression for sequence generation. This enables rapid inference leveraging KV Cache while benefiting from the performance gains of autoregressive priors. However, its macro-level autoregressive paradigm imposes a “shortsightedness,” sacrificing diffusion’s core essence—the global receptive field—and degenerating into a local model. It also introduces irreversibility similar to AR, preventing modifications to already written content.
Current research often seeks a balance between the following trade-offs: stronger global planning capabilities necessitate sacrificing inference speed, while better autoregressive priors and faster inference require sacrificing global perspective. We contend that this dichotomy is not irreconcilable. Future architectures should not be inherently AR models with minor local diffusion biases, but rather novel diffusion frameworks that preserve semi-AR advantages while retaining global vision and planning capabilities. By reintroducing global capabilities into Block Diffusion models, we aim to break the impossible triangle of speed vs. quality vs. global consistency in discrete diffusion, resulting in a unified architecture that combines the strengths of both approaches.
@@
In this study, we propose a novel generative framework called Diffusion in Diffusion (also known as Structured Block Diffusion), designing a multi-stage “draft-then-revise” paradigm to bring back the global ability to the aforementioned semi-autoregressive diffusion models. Inspired by the human writing process—rapidly drafting a first version followed by holistic revision to ensure coherence—our approach decomposes generation into progressive stages: During the initial drafting phase, the model employs a small blocksize for fast, efficient generation, focusing first on specific grammar and local content. In subsequent revision phases, the model re-examines and refines the previously generated sequence using a larger blocksize (global receptive field). We also innovatively introduce a Remasking mechanism into Block-Diffusion Models, which selectively transforms tokens within the generated sequence back into noise. This enables the model to resample these positions within a more holistic context for optimization.
@@ 4.2 Ablations and Analysis
5 Related Works
Discrete Diffusion Language Models
Discrete diffusion models have now emerged as strong competitors to autoregressive paradigms. They are renowned for their non-autoregressive parallel generation mechanism, which offers greater controllability and flexibility than traditional AR models. Early explorations like D3PM (Austin et al., 2021) modeled the forward noise process using a transition matrix that extended the Gaussian diffusion model to a discrete state space. Subsequently, masked discrete diffusion models (Gu et al., 2022; Li et al., 2022) gained prominence by explicitly incorporating a mask token to assist in reversing data corruption. Recently, approaches like SEDD (Lou et al., 2024) and MDLM (Sahoo et al., 2024) unified discrete diffusion processes with continuous time steps. By modulating noise timestep into the model, they further reduced the language modeling perplexity of diffusion models. Recent attempts to scale masked discrete diffusion models to 7B-scale (Nie et al., ; Ye et al., 2025) and beyond have demonstrated advantages in planning, mathematical, and coding tasks, further expanding their potential. The core architectural advantage of these models lies in their global bidirectional attention mechanism, which enables exceptional performance in global text structure planning and infilling tasks. However, this mechanism precludes the use of key-value caching techniques, resulting in quadratic inference complexity with sequence length. This limitation severely restricts their applicability in long-context generation tasks.
Semi-Autoregressive Block Generation
To address the efficiency challenges, researchers proposed the semi-autoregressive paradigm. Characterized by autoregressive behavior at the macro level and diffusion at the micro level, this approach leverages KV Cache to reduce computational while preserving a certain degree of local bidirectional parallelism. SSD-LM (Han et al., 2023) initially introduced the semi-AR diffusion model, achieving linear inference complexity through iterative text block generation. Subsequently, the Block Discrete Denoising Diffusion Model (Arriola et al., 2025) further optimized this paradigm. It employs inter-block autoregression and intra-block bidirectional diffusion, enabling KV cache utilization while allowing flexible control over generation length. However, this macro-level autoregressive approach also introduces inherent compromises. In semi-AR models, blocks can be abstracted as metatokens. These metatokens are constrained by strict unidirectional and irreversible dependencies, resulting in the loss of diffusion models’ inherent global planning capabilities. Moreover, early stage generation errors become irrevocable and cannot be corrected based on subsequent context, inevitably compromising consistency in long-text generation.
Iterative Refinement and Multi-Stage Generation
Our approach is also closely related to the ”Draft-then-Revise” paradigm found in multi-stage generation. In the AR domain, Speculative Decoding (Leviathan et al., 2023; Chen et al., 2023) leverages a small model to rapidly draft tokens and a large model to verify them, aiming to accelerate inference without degrading quality. More fundamentally, there are models like the Levenshtein Transformer (Gu et al., 2019) optimize text through iterative insertion and deletion operations. While Speculative Decoding mainly focuses on speeding up generation, our proposed Diffusion in Diffusion focuses on improving quality by reintroducing global planning and coherence to BD3-LMs. Rather than simply rejecting tokens based on likelihood, we use snapshot confidence to identify structural defects and bring back the global receptive field for repair. Fundamentally, our approach is a nested iterative diffusion process carefully designed to restore the long-range modeling capabilities of global diffusion within an efficient block-based diffusion framework.
6 Conclusion
This work aims to address the fundamental trade-off of global coherence and inference efficiency in semi-autoregressive diffusion models. We propose Diffusion in Diffusion, a structural diffusion framework that improves both the global coherence and the generation quality of existing block diffusion models with minimal overhead. This approach introduces a multi-stage refinement mechanism based on snapshot confidence and a mix-scale training strategy, re-injecting the long-range planning capability of Global Diffusion into an efficient block-generation framework. Results demonstrate that this Draft-then-Revise paradigm achieves state-of-the-art generation perplexity performance on OpenWebText with exceptional data efficiency, significantly narrowing the performance gap with traditional autoregressive models.